Torrid Romance Publishers has total receivables of $3040 which represent 20 days' sales.Average total assets are $73000. The firms operating profit margine is 5%. Find the firm's ROA and asset turnover ratio?
Answer #1

Total sales = $3,040 × 365/20 = $55,480

Asset turnover ratio = $55,480/$73,000 = 0.76

ROA = Asset turnover × Operating profit margin = 0.76 × 0.05 = 0.038 or 3.80%